OpenStudy (anonymous):

can someone help me with this equation i'm supposed to simplify it (3^ √6*4√ 6)^12

OpenStudy (anonymous):

see your question again.

OpenStudy (pawanyadav):

What you mean by ^ that symbol

OpenStudy (anonymous):

the 3 is and exponent so i put that sign to mean the 3 is an exponent

OpenStudy (anonymous):

am i correct? \[\left( 3^{\sqrt{6}}*4^{\sqrt{6}}\right)^{12}\]

OpenStudy (anonymous):

the 3 and 4 is supposed to be the smaller numbers and the number 6 is supposed to be big and the 12 is correct

OpenStudy (anonymous):

is it \[\left( \sqrt[3]{6} *\sqrt[4]{6}\right)^{12}\] ?

OpenStudy (anonymous):

yes thats the equation

OpenStudy (anonymous):

if so then \[=\left( 6^{\frac{ 1 }{ 3 }}*6^{\frac{ 1 }{ 4 }} \right)^{12}\] \[=\left( 6^{\frac{ 1 }{ 3 }+\frac{ 1 }{ 4 }} \right)^{12}=\left( 6^{\frac{ 4+3 }{ 12 }} \right)^{12}=6^7\]

OpenStudy (anonymous):

how did you get rid of the square root symbol

OpenStudy (pawanyadav):

That not square root that means 6 raise to the power 1/3 and 1/4

OpenStudy (anonymous):

\[\sqrt[n]{x}=x ^{\frac{ 1 }{ n }}\]

OpenStudy (anonymous):

\[\sqrt{x}=x ^{\frac{ 1 }{ 2 }}\]

OpenStudy (pawanyadav):

So when when you add 1/3 and 1/4 You will get 7/12 means 6 raised to the power 7/12 .

OpenStudy (pawanyadav):

Also there is whole power 12 So multiply 7/12 and 12 So you got 7 means the final Answer is 6^7.

OpenStudy (anonymous):

ok i think i got it thanks for helping me